# Prediction of flow and elastic stresses in a viscoelastic turbulent channel flow using convolutional neural networks

### v1

@@ Abstract
Neural-network models have been employed to predict the instantaneous flow close to the wall in a viscoelastic turbulent channel flow. The numerical simulation data at the wall is utilized to predict the instantaneous velocity fluctuations and polymeric-stress fluctuations at three different wall-normal positions. Apart from predicting the velocity fluctuations well in a hibernating flow, the neural-network models are also shown to predict the polymeric shear stress and the trace of the polymeric stresses at a given wall-normal location with reasonably good accuracy. These non-intrusive sensing models can be integrated in an experimental setting to construct the polymeric-stress field in turbulent flows, which otherwise may not be directly quantifiable in experimental measurements.
@@
Viscoelastic fluids are widely used in industrial processes, and an understanding of complex-fluid behaviour becomes crucial for enterprises working with non-Newtonian flows. Further, in real-world scenarios, turbulent flows predominate, exhibiting chaotic and multi-scale dynamics. The turbulent flows of purely viscoelastic fluids have important technological implications due to increased mixing efficiencies at low Reynolds numbers and have also piqued the interest in drag-reduction and flow control communities at high Reynolds numbers. The addition of a tiny amount of polymer (parts per million) has proven efficient in reducing friction drag in pipe flows (Virk, 1971), leading to substantial energy savings in fluid-transport applications. Drag reduction by polymers (elasticity) is related to their ability to modify coherent structures in wall-bounded turbulence (Dubief et al., 2004, 2005). Elasticity influences the turbulent cycle in two ways: by attenuating near-wall vortices, but at the same time increasing the streamwise kinetic energy of the near-wall streaks. Additionally, Xi & Graham (2010, 2012) suggested that the turbulent flow is characterized by an alternating sequence of active and hibernating phases. These phases are distinguished by flow structures exhibiting strong vortices and wavy streaks during the active phase, and weak streamwise vortices during the hibernation phase, with viscoelastic flows characterized by increased hibernation intervals. Additional insights into the influence of polymer additives on drag reduction are detailed in Xi (2019).
Identifying such effects of elasticity sparked an interest to detect and understand drag-reducing behaviour in the presence of both elasticity and plasticity in the fluid. Le Clainche et al. (2020) analysed the simulation data of Rosti et al. (2018) using high-order dynamic mode decomposition, and compared the modes in complex fluids (non-Newtonian fluids) with those in Newtonian fluids. Their results indicated that elasticity and plasticity of the complex fluids have similar effects on the coherent structures; in both cases, the flow is dominated by long streaks disrupted by rapid, localised perturbations. On the other hand, the Newtonian flow displays short streaks and more chaotic dynamics. Izbassarov et al. (2021) found that the largest amount of drag reduction is achieved with a combination of finite elasticity and plasticity, and while the highly plastic flow (high Bingham numbers) relaminarizes, elasticity affects the relaminarization in a complex and non-monotonic fashion.
However, when it comes to canonical (and practical) flows of interest, direct numerical simulations (DNSs) of viscoelastic (and/or elasto-viscoplastic) flow face challenges due to substantial computational costs associated with capturing the diverse physical mechanisms driving the flow. On the other hand, experimental investigations of drag reduction in viscoelastic flows encounter limitations stemming from near-wall measurements and the capabilities of experimental techniques to accurately quantify the flow, without perturbing it. While a complete description of viscoelastic turbulence would require characterization of both velocity and polymeric stresses. However accessing such polymer deformation directly from experimental measurements remains a challenging goal.
@@
In recent years, machine-learning methods have provided a significant advance over prior methodologies in various applications within the field of fluid mechanics (Vinuesa, 2024). Specifically in the domain of viscoelastic flows, researchers have explored the predictability of polymer stress components from velocity gradient using neural networks (Nagamachi & Tsukahara, 2019). Towards estimation of fluid flow, neural-network models have demonstrated excellent capabilities in predicting the instantaneous state of the Newtonian turbulent flow using quantities measured at the wall (Guastoni et al., 2021). Hence in the present study, the idea of non-intrusive sensing has been applied to viscoelastic turbulent channel flow to predict the velocity fluctuations and polymeric stress components near the wall using the quantities measured at the wall. To this end, the convolutional-neural-network (CNN) models are employed to predict the two-dimensional velocity fluctuations and polymeric-stress fluctuation fields at different wall-normal distances. The present work highlights the capability of a data-driven approach to model turbulence in complex-fluids flows. In addition, the developed non-intrusive sensing models will also find useful applications in experimental settings and in closed-loop control of wall-bounded turbulence in viscoelastic flows.
@@ 2.1 Dataset
The viscoelastic turbulent channel flow is simulated using a finite-difference-based in-house code on a computational domain of size in the streamwise , wall-normal and spanwise directions , respectively as shown in figure 1. The readers are referred to Izbassarov et al. (2021) for a complete description of the viscoelastic turbulent channel simulation employed in this study. The computational domain is uniformly discretized using grid points along and , respectively. The spatial resolution of is chosen to fully resolve the turbulent scales in the viscoelastic turbulent flow (Rosti et al., 2018). Here, the superscript ‘+’ denotes the scaling in terms of friction velocity (, where corresponds to the wall-shear stress) as the velocity scale and viscous length as the length scale. Note that the value of obtained with is lower than that in the Newtonian case due to skin-friction reduction. Variation of the averaged wall-shear rate obtained with is compared against the Newtonian case in figure 2. Here, corresponds to the wall-normal derivative of the streamwise velocity and denotes the spatial averaging in and directions in the channel. The Reynolds decomposition of an instantaneous field is denoted as , with identifying the fluctuations. From figure 2, identifying the hibernation intervals (regions with low wall-shear stress) using area-averaged wall-shear rate as performed in Xi & Graham (2010) with a threshold corresponding to of the mean shear-rate, we observe the presence of such low-drag events at , where polymer stretching is weakest. Note that the choice of threshold is arbitrary here and a definitive choice of the threshold value is absent in the literature. Effectively, for the considered viscoelastic turbulent flow at , we observe a drag reduction of roughly for the set of considered parameters in this study. From figure 2, it is evident that the wall-fields (which are provided as inputs to FCN, see figure 1, §2.2) are significantly varying from the statistical mean for a considerable fraction of the total time. Thus, in this work we aim to build a neural-network model that can predict viscoelastic turbulence quantities of interest, not only in the mean-flow but also in extreme wall-shear events with particular interest in hibernation intervals.
To this end, a database consisting of instantaneous fluctuation fields of wall-shear-stress components, wall-pressure fluctuation, two-dimensional velocity-fluctuation and polymeric-stress-fluctuation fields are obtained at different wall-normal locations, and ( and , respectively). Note that these wall-normal locations correspond respectively to and for a Newtonian turbulent channel flow and hence, for simplicity we refer to these locations as and , respectively in this study. The simulations are run for time-units and a total of 40,600 samples is obtained with a sampling period of for training the network model. The sampled instantaneous two-dimensional fields are down-sampled to a resolution of in and , respectively. The fields at both the walls are utilized in this study, and they are split into training and validation set with a ratio of 4 to 1. The network models are evaluated with the samples in the test dataset which consists of 10,000 samples. The samples in the test dataset are chosen from a time-interval (in the sampled time series) that corresponds to at least 60 flow-through times apart from the samples in the training dataset to ensure minimal autocorrelation between the samples in training and test dataset. The number of samples in the test dataset is chosen to obtain convergence of second-order turbulence statistics.
@@ 2.2 Neural-network model
In this work, a fully convolutional neural network (FCN) similar to the one proposed by Guastoni et al. (2021) is used, with increased hidden layers (see figure 1) to obtain a more complex combination of abstract turbulent features identified by the kernels in the network. Here, we utilize an existing architecture, acknowledging that further enhancements could be achieved with newer architectures that require extensive datasets. Our focus is on proposing a methodology for viscoelastic stress predictions in turbulent flows aimed towards experimental applications and in establishing baseline performance using current convolutional architectures. The considered FCN consists of 30 hidden layers with a total number of trainable parameters amounting to . The convolution operations are followed by a batch normalization and a rectified-linear-unit (ReLU) activation function. The inputs to the network are normalized using respective mean and standard deviation of the fields from the training dataset and the outputs are normalized using the corresponding standard deviation values. The choice of loss function in the network is the mean-squared error (MSE) between the instantaneous predicted and DNS fields, which helps the network to learn the large-scale features first and then progressively optimize the trainable parameters to minimize the errors at finer scales (Xu et al., 2019).
@@
In this study, three types of predictions have been undertaken to highlight the capability of FCN models to reconstruct the near-wall visco-elastic turbulence fields. In V-predictions (indicating velocity predictions), the streamwise wall-shear, spanwise wall-shear and pressure field at the wall are utilised to predict the streamwise, spanwise and wall-normal velocity-fluctuations. This allows us to assess whether velocity fields can also be predicted in a viscoelastic turbulence exhibiting periods of hibernation. The performance of neural-network models in predicting the fluctuations of polymeric-shear stress () and fluctuations of trace of polymer stress () at a given wall-normal location using the true velocity-fluctuation fields at the same location, and they are denoted as E-predictions (signifying prediction of elastic stress quantities of interest). Finally in V-E-predictions, the FCN model is used to predict the fluctuations of polymeric-shear stress and diagonal components of polymer-stress tensor at a target wall-normal distance directly from wall inputs, with auxillary predictions of corresponding velocity-fluctuations at the considered wall-normal location. 111The auxiliary predictions of velocity-fluctuations at a wall-normal location is utilized in V-E predictions to obtain an increase in the accuracy of prediction of polymeric stress quantities.
@@
The network performance is evaluated from a statistical point of view in terms of the relative-error in predicting the corresponding root-mean-squared (RMS) quantities between the true (DNS) fields from test dataset and predicted fields from FCN (indicated by ). The mean-absolute error between the predictions and DNS fields (denoted by ) is also reported for different types of predictions. Note that the performance metrics reported in this study are obtained from the mean of at least three different network models to include the effects of stochasticity introduced by the random initialization of kernel weights in FCN and random sampling of mini-batches during the training process. The instantaneous correlation coefficient between the predicted and DNS fields, averaged over the samples in the test dataset, is also highlighted. To evaluate the distribution of energy in different length scales, a comparison of the pre-multiplied two-dimensional (2D) power-spectral density (PSD) between DNS fields and the predictions is performed. Here, is the power-spectral density obtained for the quantity ‘’ and respectively denote the wavenumbers in streamwise and spanwise directions with the corresponding wavelengths in viscous units denoted by and .

### v2

@@ Abstract
Neural-network models have been employed to predict the instantaneous flow close to the wall in a viscoelastic turbulent channel flow. Numerical simulation data at the wall is utilized to predict the instantaneous velocity-fluctuations and polymeric-stress-fluctuations at three different wall-normal positions in the buffer region. The ability of non-intrusive predictions has not been previously investigated in non-Newtonian turbulence. Our analysis shows that velocity-fluctuations are predicted well from wall measurements in viscoelastic turbulence. The models exhibit enhanced accuracy in predicting quantities of interest during the hibernation intervals, facilitating a deeper understanding of the underlying physics during low-drag events. The neural-network models also demonstrate a reasonably good accuracy in predicting polymeric-shear stress and the trace of the polymer stress at a given wall-normal location. This method could be used in flow control or when only wall information is available from experiments (for example, in opaque fluids). More importantly, only velocity and pressure information can be measured experimentally, while polymeric elongation and orientation cannot be directly measured despite their importance for turbulent dynamics. We therefore study the possibility to reconstruct the polymeric-stress fields from velocity or pressure measurements in viscoelastic turbulent flows. The results are promising but also underline that a lack of small scales in the input velocity fields can alter the rate of energy transfer from flow to polymers, affecting the prediction of the polymer-stress fluctuations. The present approach not only aids in extracting polymeric-stress information but also gives information about the link between polymeric-stress and velocity fields in viscoelastic turbulence.
@@
Viscoelastic fluids are widely used in industrial processes, and an understanding of complex-fluid behaviour becomes crucial for enterprises working with non-Newtonian flows. Further, in real-world scenarios, turbulent flows predominate, exhibiting chaotic and multi-scale dynamics. The turbulent flows of purely viscoelastic fluids have important technological implications due to increased mixing efficiencies at low Reynolds numbers and have also piqued the interest in drag-reduction and flow control communities at high Reynolds numbers. The addition of a tiny amount of polymer (parts per million) has proven efficient in reducing friction drag in pipe flows (Virk, 1971), leading to substantial energy savings in fluid-transport applications. Such observation sparked intense research into the interaction between flow dynamics and polymers, in the context of skin-friction reduction. Lumley (1969) postulated that the drag reduction attributed to polymer molecules arises from the extension of polymers, thereby increasing the effective extensional viscosity of the solution. This increase in extensional viscosity leads to the damping of small eddies, resulting in increased buffer-layer thickness. However, De Gennes (1986) attributed the drag reduction to elastic effects rather than the viscous properties of polymers. This theory assumes that the turbulent kinetic energy (TKE) is stored as elastic energy by the polymers, thereby influencing the energy cascade. Nevertheless, neither theory offers a comprehensive description of polymer induced drag reduction.
Various experimental investigations in channel and pipe flows (Pinho & Whitelaw, 1990; Wei & Willmarth, 1992; Den Toonder et al., 1997; Warholic et al., 1999, 2001; Ptasinski et al., 2001, 2003) have shown that polymers induce changes in the turbulence structure rather than simply attenuating it. These changes are characterized by an increase in streamwise turbulence intensity alongside a decrease in wall-normal turbulence. With access to direct numerical simulation (DNS) techniques (Sureshkumar et al., 1997), a deeper insight into the intricate interaction between polymers and flow fields has been probed. Notably, Dubief et al. (2004, 2005) pointed out that the polymers exhibited a structured energy exchange with the flow, primarily occurring around near-wall vortices. This increased polymer activity in the buffer region suggests that polymers are entrained around the vortices, resulting in the torque due to polymer stress opposing the rotation of the streamwise vortices (Kim et al., 2007), thereby extracting energy from near-wall vortices. Consequently, polymer stretching weakens the near-wall coherent structures, leading to a reduction of skin friction (Stone et al., 2002; Dubief et al., 2004; Li & Graham, 2007; Kim et al., 2007, 2008). Besides damping near-wall vortices, polymers also enhance the streamwise kinetic energy in near-wall streaks, thereby the net balance of these gives rise to a self-sustained drag-reduced turbulent flow (Dubief et al., 2004, 2005). Additionally, Xi & Graham (2010, 2012) suggested that the turbulent flow is characterized by an alternating sequence of active and hibernating phases. These phases are distinguished by flow structures exhibiting strong vortices and wavy streaks during the active phase, and weak streamwise vortices during the hibernation phase, with viscoelastic flows characterized by increased hibernation intervals. Additional insights into the influence of polymer additives on drag reduction are detailed in Xi (2019).
Overall, the studies indicate that drag reduction due to polymer molecules results from the dynamic interaction between polymer molecules and turbulence. Nonetheless, the study of polymer drag reduction is also limited by the fact that the exact behaviour of polymers in turbulent flows is not fully deciphered. Moreover, the intermittent dynamics of polymers near the wall are predominantly explored through numerical simulations rather than experimental investigation of polymer stress. Consequently, there exists a scope to gain insights into polymer stress through potential experimental measurements of flow fields. However, experimental investigations of drag reduction in viscoelastic flows encounter limitations stemming from near-wall measurements and the capabilities of experimental techniques to accurately quantify the flow, without perturbing it. While a complete description of viscoelastic turbulence would require characterization of both velocity and polymeric stresses, accessing such polymer deformation directly from experimental measurements remains a challenging goal (Stone & Graham, 2003).
The phenomenon of drag reduction in viscoelastic turbulent flows has been explored from theoretical, experimental and numerical perspectives. However, in recent years machine-learning methods have expanded the possibilities in simulating (Raissi et al., 2020; Eivazi et al., 2022; Vinuesa & Brunton, 2022), predicting (Guastoni et al., 2021; Eivazi et al., 2021; Yousif et al., 2023a, b) and controlling (Vignon et al., 2023; Guastoni et al., 2023) wall-bounded Newtonian turbulent flows (Vinuesa, 2024). Data-driven methods also hold the potential for enhancing the understanding of the role of coherent structures in Newtonian (Cremades et al., 2024) and viscoelastic turbulence dynamics (Le Clainche et al., 2020). In the domain of elastic stress predictions, researchers have explored the predictability of polymer stress components from velocity gradient using neural networks (Nagamachi & Tsukahara, 2019) and have also developed a framework for rheological model discovery from velocity field and information corresponding to boundary conditions and initial condition of polymer stress (Thakur et al., 2024). These examples underscore the possibilities offered by data-driven methods to advance the understanding of polymers and flow field interaction in viscoelastic turbulent flow. However, a critical first step is to estimate the instantaneous polymer behaviour in an experimental setting. But, experimental studies encounter difficulties in accurately quantifying macromolecular extension, thereby significantly restricting the ability to fully characterize polymer behaviour (Stone & Graham, 2003). Moreover, near-wall measurements of flow fields in turbulent channel flows can also pose challenges. Nevertheless, non-linear machine-learning tools offer possibilities for improving experimental flow measurements and estimating flow and scalar quantities (Vinuesa et al., 2023; Eivazi et al., 2024). Neural-network models, and in particular fully-convolutional network (FCN) models have demonstrated excellent capabilities in predicting the instantaneous state of the flow using quantities measured at the wall in Newtonian turbulent flow (Guastoni et al., 2021).
@@
Hence in the present study, the idea of non-intrusive sensing has been applied to viscoelastic turbulent channel flow to predict the velocity fluctuations and polymeric-stress components near the wall using the quantities measured at the wall. To this end, FCN models are employed to predict the two-dimensional velocity fluctuations and polymeric-stress fluctuation fields at different wall-normal distances. The present work highlights the capability of a data-driven approach to model turbulence in complex-fluids flows. As a first step, directly providing a possible experimental measurement of wall-quantities or near-wall velocity fluctuations as inputs to the network model allows for an estimation of the polymer stress quantities in the near-wall region from experimental observations. Furthermore, the developed non-intrusive sensing models will also find useful applications in closed-loop control of wall-bounded turbulence in viscoelastic flows.
The paper is organised as follows. The methodology employed in the present work is introduced in §2. A description on the dataset to train the network models is provided in §2.1 and an overview of the network models can be found in §2.2. A filter-based approach to identify the effects of small-scales on FCN predictions is detailed in §2.3. In §3, the results obtained with different network models are discussed. The performance of the network models in different prediction types is outlined in §3.1–§3.3. The effects of filtering the small scales in the input velocity fluctuations on the predictions by FCN is discussed in §3.4–§3.5 and an interpretation of the predictions by FCN is provided in §3.6. Additional details regarding the error metrics for different types of predictions and the analysis of predictions of polymer shear stress can be respectively found in Appendix A–C.
@@ 2.1 Dataset
In order to estimate the turbulent dynamics of the viscoelastic fluid from velocity fields, direct numerical simulation of viscoelastic turbulent channel flow is performed to generate the data necessary for modelling the relationship between near-wall velocity-fluctuations and polymer-stress-fluctuation quantities of interest using a fully-convolutional neural-network model. For the direct numerical simulation, a macroscopic continuum description of fluid-polymer interaction is considered. Specifically, a homogeneous dilute mixture of polymer in the solvent fluid is modelled with the finitely extensible non-linear elastic with Peterlin closure (FENE-P) constitutive relation (refer equation (4)). Although more sophisticated models (Watanabe & Gotoh, 2014; Shen et al., 2022) could be considered, Stone & Graham (2003) demonstrated that the FENE-P model can still yield reasonable predictions of the spatial distribution of stresses due to polymers and indicated that while the transient extension for the FENE-P model compares favourably to more detailed model such as bead-spring-chain models, it may overestimate elastic stresses in turbulent flow. Nonetheless, the FENE-P model can reproduce features of polymer drag reduction, including the maximum-drag reduction asymptote (Graham, 2014). Hence, in the present study, FENE-P model is utilized with the notion that it accurately describes the polymer physics of the considered viscoelastic liquid.
In the macroscopic description of the polymer model, we consider the statistics of end-to-end vector to characterize the average of the end-to-end vectors of a large number of polymer macro-molecules contained in a fluid particle. An essential statistical quantity in this description is the second-order correlation of the orientation vector , known as the conformation tensor which is obtained by:
where corresponds to the statistical average over a fluid particle and corresponds to the root mean square (RMS) of all the end-to-end vectors inside a fluid particle at rest.
The dimensional incompressible Navier–Stokes equations coupled with the evolution equation for the polymer conformation tensor are given by:
where is the velocity with corresponding components in streamwise, wall-normal and spanwise directions represented by and is the pressure. In the remainder of the text, mean velocity components are denoted by and the corresponding fluctuations by . In equation (2), is the density of the fluid with denoting the time-scale of flow and denoting the relaxation time-scale of polymer stress. The solvent and polymer viscosity of the fluid are given by , respectively and they depend on the polymer concentration. The Peterlin function accounts for the finite length of the polymer molecules with denoting the upper limit of the normalized polymer extension length, the point after which, the polymers cannot absorb more energy from the flow.
The polymer stress can be retrieved from the conformation tensor using the Kramer relationship:
@@
The dataset for training and evaluation of the network model is obtained through a DNS of turbulent channel flow of viscoelastic fluid at a Reynolds number based on the bulk velocity of ( denotes the total kinematic viscosity of the fluid), which corresponds to a friction Reynolds number (where is defined in terms of and friction velocity ) for a Newtonian fluid. In this study, the turbulent channel flow simulations are performed at a Weissenberg number . The ratio of polymeric viscosity to the total viscosity , which is denoted by , is set to . The maximum polymer extensibility is set to . The difficulties associated with proper rheological characterization of real fluids by adequate constitutive equations is an important area of research on its own and rather, we assume that the adopted model adequately describes the intended fluid properties.
@@
The viscoelastic turbulent channel flow is simulated using a finite-difference-based in-house code on a computational domain of size in the streamwise , wall-normal and spanwise directions , respectively as shown in figure 1. The readers are referred to Izbassarov et al. (2021) for a complete description of the viscoelastic turbulent channel simulation employed in this study. The computational domain is uniformly discretized using grid points along and , respectively. A spatial resolution of is chosen to fully resolve the turbulent scales in the viscoelastic turbulent flow (Rosti et al., 2018). Here, the superscript ‘+’ denotes the scaling in terms of the friction velocity (, where corresponds to the wall-shear stress) and viscous length . Note that the value of obtained with is lower than that in the Newtonian case due to skin-friction reduction. Variation of the averaged wall-shear rate obtained with is compared against the Newtonian case in figure 2. Here, corresponds to the wall-normal derivative of the streamwise velocity and denotes the spatial averaging in and directions in the channel. From figure 2, identifying the hibernation intervals (regions with low wall-shear stress) using area-averaged wall-shear rate as performed in Xi & Graham (2010) with a threshold corresponding to of the mean shear-rate, we observe the presence of such low-drag events at . Note that the choice of threshold is arbitrary here and a definitive choice of the threshold value is absent in the literature. Effectively, for the considered viscoelastic turbulent flow at , we observe a drag reduction of roughly for the set of considered parameters in this study. From figure 2, it is evident that the fields at the wall (which are provided as inputs to FCN, see figure 1, §2.2) significantly deviate from the statistical mean for a considerable fraction of the total time. Thus, in this work, we aim to build a neural-network model that can predict viscoelastic turbulence quantities of interest, not only in the mean-flow but also in extreme wall-shear events with particular interest in hibernation intervals.
To this end, a database consisting of instantaneous fluctuation fields of wall-shear-stress components and wall-pressure fluctuation, as well as the two-dimensional velocity-fluctuation and polymeric-stress-fluctuation fields obtained at different wall-normal locations, and ( and , respectively) is generated. Note that these wall-normal locations correspond respectively to and for a Newtonian turbulent channel flow and hence, for simplicity, we refer to these locations as and , respectively in this study. Buffer region is probed because of its importance in terms of production and dissipation of turbulent energy and is significantly modified in drag-reducing flows with polymer additives (Den Toonder et al., 1997; Stone & Graham, 2003).
The simulations are run for time-units and a total of 40,600 samples is obtained with a sampling period of for training the network model. The sampled instantaneous two-dimensional fields are down-sampled to a resolution of in and , respectively. The fields on both walls are utilized in this study, and they are split into training and validation set with a ratio of 4 to 1. The computational resources required to carry out the simulation and generate the necessary training dataset amounted to approximately 1.5 million core-hours.
The network models are evaluated with the samples in the test dataset which consists of 10,000 samples. The samples in the test dataset are chosen from a time-interval (in the sampled time series) that corresponds to at least 60 flow-through times apart from the samples in the training dataset to ensure minimal auto-correlation between the samples in the training and test dataset. The temporal length for the number of samples considered in the test dataset corresponds to time units, whereas a temporal averaging over 10–15 eddy-turnover times is sufficient to obtain good convergence of turbulence statistics (Li et al., 2006).
@@ 2.2 Neural-network model
In this work, a fully convolutional neural network similar to the one proposed by Guastoni et al. (2021) is used, with an increased number of hidden layers (see figure 1) to obtain a more complex combination of abstract turbulent features identified by the kernels in the network. Here, we utilize an existing architecture, acknowledging that further enhancements could be achieved with newer architectures that require extensive datasets. Our focus is on proposing a methodology for viscoelastic stress predictions in turbulent flows aimed towards experimental applications and in establishing baseline performance using current convolutional architectures. The considered FCN consists of 30 hidden layers with a total number of trainable parameters amounting to . The convolution operations are followed by batch normalization and a rectified-linear-unit (ReLU) activation function. The inputs to the network are normalized using respective mean and standard deviation of the fields from the training dataset and the outputs are normalized using the corresponding standard deviation values. The choice of loss function () in the network is the mean-squared error (MSE) between the instantaneous predicted and DNS fields:
which helps the network to learn the large-scale features first and then progressively optimize the trainable parameters to minimize the errors at finer scales (Xu et al., 2019). The subscripts ‘’ and ‘’ are respectively used to denote the DNS samples and corresponding predictions from FCN. Each network model is trained using graphics-processing units (GPUs), amounting to approximately 3,500 GPU-hours for training a single network model.
@@
In this study, three types of predictions have been undertaken to highlight the capability of FCN models to reconstruct the near-wall visco-elastic turbulence fields. In V-predictions (indicating velocity predictions), the streamwise wall-shear, spanwise wall-shear and pressure field at the wall are utilised to predict the streamwise, spanwise and wall-normal velocity-fluctuations. This allows us to assess whether velocity fields can also be predicted in viscoelastic turbulence exhibiting periods of hibernation. The performance of neural-network models in predicting the fluctuations of polymeric-shear stress () and fluctuations of trace of polymer stress () at a given wall-normal location using the true velocity-fluctuation fields at the same location are denoted as E-predictions (signifying prediction of elastic stress quantities of interest). Finally in V-E-predictions, the FCN model is used to predict the fluctuations of polymeric-shear stress and diagonal components of polymer-stress tensor at a target wall-normal distance directly from wall inputs, with auxiliary predictions of corresponding velocity-fluctuations at the considered wall-normal location 111The auxiliary predictions of velocity-fluctuations at a wall-normal location is utilized in V-E predictions to obtain an increase in the accuracy of prediction of polymeric-stress quantities.. It is worth noting that the mean of polymer shear stress can be obtained either from experimentally quantified stress-deficit (Warholic et al., 1999) or from numerical simulations. Additionally, the FCN model performs well in predicting the mean of polymer-stress quantities, which stems from the use of the mean-squared error as a loss function that tends to regress to the mean and hence the relative errors in predicting the mean polymer stress quantities from FCN are lower than . Consequently, our primary focus in this study is on retrieving the instantaneous fluctuations of polymer-stress using corresponding fluctuations of input quantities.
The mean-absolute error between the predictions and DNS fields (denoted by ) is reported for different types of predictions, which is calculated as
The error metrics described in this section is computed for each component of and for the polymer-stress quantities of interest (refer Appendix A). The network performance is also evaluated from a statistical point of view in terms of the relative error in predicting the corresponding root-mean-squared (RMS) quantities between the true (DNS) fields from test dataset and the fields predicted by the FCN (indicated by ), and is given by:
whereas the instantaneous correlation coefficient between the predicted and the DNS fields is defined as:
@@
with corresponding to the average in space () and time (; denotes the samples in the test dataset) and subscript refers to root-mean-squared quantities. Note that the performance metrics reported in this study are obtained from the mean of at least three different network models to include the effects of stochasticity introduced by the random initialization of kernel weights in FCN and random sampling of mini-batches during the training process. The instantaneous correlation coefficient between the predicted and DNS fields averaged over the samples in the test dataset, is also highlighted. To evaluate the distribution of energy across the various scales in the flow, a comparison of the pre-multiplied two-dimensional (2D) power-spectral density (PSD) between DNS fields and the predictions is performed. Here, is the power-spectral density obtained for the quantity ‘’ and respectively denote the wavenumbers in streamwise and spanwise directions, with the corresponding wavelengths in viscous units denoted by and .
2.3 Low-pass filtering of velocity fluctuations
In order to identify the effects of small-scale features in the input velocity-fluctuation fields on the predictions of polymer stress (see §3.5–§3.6) and thereby to recommend scale requirements for inputs from possible experimental investigations, we employ low-pass filtering of velocity-fluctuations with a threshold wavelength denoted by . This threshold represents the smallest wavelength of the turbulent scales present in the input velocity-fluctuation fields to the FCN. For instance, in velocity-fluctuation fields sampled from DNS with data-points respectively in the streamwise and spanwise directions, the minimum wavelengths correspond to and hence, for the DNS sampled fields. By filtering the sampled DNS velocity-fluctuation fields at a certain to contain small-scales above , we ascertain that . However, the predictions of polymer stress quantities from the FCN aim to capture all the scales above 2.5 viscous-length units, regardless of the input . In other words, the sampled DNS polymer-stress quantities serve as the true reference for the supervised training process of FCN for all the different network models with corresponding of input velocity-fluctuations.
The filtering is performed in the wavenumber space and the corresponding procedure for the filtering process is outlined in Algorithm 1.
The relative decrease in the turbulent kinetic energy in the input velocity-fluctuations sampled from DNS (denoted by ) and after the filtering process (denoted by ) for a given is quantified as:
where corresponds to the spatial and temporal averaging of data yielding a scalar output. Here, corresponds to the sequence of sampled velocity-fluctuation snapshots from DNS and is the sequence of respective filtered fields obtained from filtering process.